OpenStudy (anonymous):

Simplify the expression. 34 + 12 ÷ 3 - (1-9) a.)75 b.)21 c.)93 d.)39

myininaya (myininaya):

order operations tells us to do whats inside grouping symbols first

myininaya (myininaya):

then do exponents but here you dont have any.

myininaya (myininaya):

now do multiplication and division

myininaya (myininaya):

after that addition and subtraction

OpenStudy (anonymous):

i dont qet it

OpenStudy (anonymous):

do the multiplication or division first before the substraction or addition

OpenStudy (anonymous):

None of your listed answers are correct. First operation (1-9)=-8 now you have 34+12/3-(-8) second operation -(-8)=+8 34+12/3+8 next operation 12/3=4 34+4+8=46 there is a mnemonic trick. Please excuse my dear aunt sally Parentheses firs, exponentiation next, multiplication or division next, addtion or subtraction last. I hope this is the answer you are looking for.